Galvanometer together with an unknown resistance in series is connected across two identical batteries each of 1.5 holes in the batteries are connected in series the Galvanometer records a current of 1 amps and when the batteries are connected in parallel the current is 6 amps what is internal resistance of each battery

Grade:12th pass

1 Answers

Vikas TU
6 years ago
The net resistance = R+2r
Potential Difference(V) = 3V
As from the ohms law
3 = 1(R+2r)---------- 1.)
The net resistance in second case = R+r/2
V = 1.5V I =6A
1.5 = 6(R+r/2)
¼ =R+r/2 ---------(2.)
3r/2 = 11/4
r = 22/12
r =1.83 ohm